1/3

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,

Le loup-garou Varulven The Banshee (An Approach)

The Werewolf

tradukita de R. Platteau  tradukita de Sven Collberg  tradukita de Max Knight  tradukita de Alexander Gross

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle

! ”

En varulv lopp från fru m.m. en vacker afton samt begav sig till en bymagisters grav och bad: “Var snäll och konju-

gera.

One night, a banshee slunk away from mate and child, and in the

gloom went to a village teacher’s tomb, requesting him: “Inflect me, pray.”

A Werewolf, troubled by his name, Left wife and brood one night and came To a hidden graveyard to enlist The aid of a long-dead philologist.

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle

! ” mig!” Byskolläraren stod upp på blecknamnsskyltens

mässingknopp med svar till den, som satt på

pass med tåligt korslagd tass på

tass.

The village teacher climbed up

straight upon his grave stone with its plate and to the apparition said who meekly knelt before the dead:

“Oh sage, wake up, please don’t berate

me,” He howled sadly, “Just conjugate me.” The seer arose a bit unsteady Yawned twice, wheezed once, and then

was ready.

“ - On peut à ton sujet puisque nous y som-

mes Se demander encor bien des choses en som-

me. Loup-garoù ? mais avant Il faut se dire : loup-garquand ?

“Jag var-ulv 1:a pers. i sing., du, han, hon, den, det var-ulv.

Så vi voro-ulv, I voren. På de voro-ulv kom ingenting.”

“The banSHEE, in the subject’s

place; the banHERS, the possessive ca-

se. The banHER, next, is what they

call objective case–and that is all.”

“Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present ten-

se, And ‘Iswolf,’ ‘Arewolf’ in this same sen-

se.”

…  …  …  …

2/3

Loup-garcomment ? et puis Bien évidemment loup-garqui ? ” Toutes ces questions semblèrent au loup bi-

en absconses, Car en fait ce qu’il voulait c’était les

réponses.

Varulven myste gladåt hur pers. hade följts av sing. och

plur. “Fast”, bad han, “kan det inte

ges ens en enda pers. i tempus pre-

sens?”

The banshee marveled at the ca-

ses and writhed with pleasure, ma-

king faces, but said: “You did not add, so far, the plural to the singular!”

“I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied. The scholar paused–again he tried:

“ - Point de réponses, dit le puits de savoir, Etre imaginaire, nul ne peut te voir ! ” A ces mots l’animal éclata tout en larmes. “ - Oh!, même les baisers de ma douce moitié Bekänna måste vår magister, att därvidlag fanns vissa bri-

ster. En varulv sine imperfecto var contradictio in adjecto.

The teacher, though, admitted

then that this was not within his ken. “While bans are frequent”, he ad-

vised, “A she cannot be plurized.”

“A ‘Will-be-wolf?’ It’s just too long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Ut-

terly wrong! Such words are wounds beyond all

suture– I’m sorry, but you have no future.”

Ne pourrons plus me faire me sentir tout

entier. Rester un loup-garou est un sort qui

m’alarme. ” Mais comme le loup était bien poli Il remercia et puis repartit.

Varulven satt med tårad lins. Han visste att han fanns och

finns. Därtill kom också fru m.m. som även måste existera.

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned crea-

ture, said humbly “Thanks” and left the

teacher.

The Werewolf knew better–his sons

still slept At home, and homewards now he crept, Happy, humble, without apology For such folly of philology.

…  …  …  …

3/3

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Francan de R. Platteau.

Arg-825-1655 (2013-02-13 21:56:38)

Tiu ĉi francigo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Svedan de Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .